Solve the following problem in Thermodynamics: Car
Answer #1

1)

carnot efficiency = 1 - Tsink / Tsource = 1 - 300/2000 = 85%

Work output / Power input = 85%

work output = 0.85 x 500kW

= 425 kW

2) actual work output = 300kW

actual efficiency = 300 /500 = 60%

3)enegy loss = theoretical work output - actual work output = 425 - 300 = 125kW